In a journal, an article reported that the results of a peer tutoring program to help special children with disabilities learn to read. In the experiment, the children were randomly divided in two groups: the experimental group received peer tutoring along with regular instruction; and the control group received regular instruction with no peer tutoring. There were 30 children in each group. A test was given to both groups before instruction began. For experimental group, the mean score on the test was �𝑥�1� = 344.5 with sample standard deviation s1 = 49.1. For the control group, the mean score on the same test was �𝑥�2 = 354.2 with sample standard deviation s2 = 50.9. Use a 5% level of significance to test the hypothesis that there was no difference in the test of two groups before the instruction began.


1
Expert's answer
2021-12-14T13:53:42-0500

The following null and alternative hypotheses need to be tested:

"H_0:\mu_1=\mu_2"

"H_1:\mu_1\not=\mu_2"

This corresponds to a two-tailed test, for which a t-test for two population means, with two independent samples, with unknown population standard deviations will be used.

Testing for Equality of Variances

A F-test is used to test for the equality of variances. The following F-ratio is obtained:


"F=\dfrac{s_1^2}{s_2^2}=\dfrac{49.1^2}{50.9^2}=0.9305"

The critical values for "\alpha=0.05, df_1=30-1=29" degrees of freedom, "df_2=30-1=29" degrees of freedom are "F_L=0.476,F_U=2.101," then the null hypothesis of equal variances is not rejected.

Based on the information provided, the significance level is "\alpha = 0.05," and the degrees of freedom are "df=n_1-1+n_2-1=58."

Hence, it is found that the critical value for this two-tailed test is for "\alpha = 0.05, df=58" are of freedom is "t_c = 2.001717."

The rejection region for this two-tailed test is "R = \{t: |t| > 2.001717\}."

Since it is assumed that the population variances are equal, the t-statistic is computed as follows:


"t=\dfrac{\bar{X_1}-\bar{X_2}}{\sqrt{\dfrac{(n_1-1)s_1^2+(n_2-1)s_2^2}{n_1+n_2-2}(\dfrac{1}{n_1}+\dfrac{1}{n_2})}}"

"=\dfrac{344.5-354.2}{\sqrt{\dfrac{(30-1)49.1^2+(30-1)50.9^2}{30+30-2}(\dfrac{1}{30}+\dfrac{1}{30})}}"


"\approx-0.751237"

Since it is observed that "|t|=0.751237<2.001717,"  it is then concluded that the null hypothesis is not rejected.

Using the P-value approach: The p-value for two-tailed "df=58" degrees of freedom, "t=-0.751237" is "p = 0.455547," and since "p = 0.455547"

">0.05=\alpha," it is concluded that the null hypothes is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu_1" is different than "\mu_2," at the "\alpha = 0.05" significance level.